# organization=Instituto Federal de Educação Ciência e Tecnologia do Ceará (IFCE), Campus Tianguá,addressline=Av Tabelião Luiz Nogueira, s/n - Santo Antônio, city=Tianguá, postcode=62324-075, state=Ceará, country=Brazil

### v3

@@ Abstract
In this letter, we propose a relationship between the so-called Hubble-Lemaître constant $H_{0}$ and holographic complexity related to the emergence of spacetime in quantum gravity. Such a result can represent an important step to understanding the Hubble tension by introducing a quantum gravity perspective for cosmological observations: regarding the degree of quantum complexity we measure around us.
journal: Physics Letters B
@@ 1 Introduction
In recent years, a contention has arisen in cosmology, the so-called Hubble tension [1, 2, 3]. Such a problem has its origin in a controversy between early-universe measurements of the so-called Hubble-Lemaître constant $H_{0}$ provided by the Planck collaboration [4], and late-time observations provided by SHOES [5]. Such a discrepancy has been confirmed by several other late-time observations [5, 6, 7, 8, 9, 10, 11, 12], which have shown a clear distinction from the value provided by the Planck satellite.
The early-universe result is based on the analysis of the Cosmic Microwave Background radiation (CMB) which was formed when the universe was in its childhood. On the other hand, the late-time universe measurements are based on several observations related to objects formed in epochs when the universe was more old and complex such as cepheids, quasars, red giants stars, etc.
Putting it in numbers, the early universe result has provided a value of $H_{0}=67.4\pm 0.5\;km\;s^{-1}Mpc^{-1}$. On the other hand, the late-time universe measurements have provided values for $H_{0}\geq 70$. For example, the latest SHOES measurement gives $H_{0}=73.2\pm 1.3\;km\;s^{-1}Mpc^{-1}$ [12].
@@
Several proposals have been suggested to mend the $\Lambda$CDM model and explain the different values measured to $H_{0}$. An extensive list of such proposals can be found in [2]. In front of these efforts, it has been argued that a deeper breakdown of our understanding of the cosmos becomes necessary to solve the Hubble tension.
A forthcoming quantum version of cosmology, based on a still unknown theory of quantum gravity, can promote such a paradigm shift, by reformulating our view about the nature of spacetime itself. However, so far, no satisfactory proposal to solve the Hubble tension has been introduced by any candidate theory to explain the quantum aspect of the gravitational phenomena.
In the present letter, we propose a possible explanation for the origin of the Hubble tension by considering new results on quantum gravity introduced recently in [20, 21]. In such a context, we shall find out that $H_{0}$ can be related to the quantum complexity associated with the emergence of spacetime.
Consequently, the discrepancies among the $H_{0}$ values, when one considers the early and late-time measurements, can rely on a natural aspect of the cosmos: the evolution of its complexity throughout time according to the second law of holographic complexity [22]. Such a result may not only shed light on the Hubble tension but also introduce a new perspective for cosmological observations: in terms of holographic quantum complexity we measure around us.
The paper is organized as follows: in sections (2) and (3), we review the main aspects of the theory introduced in [20, 21]. In section (4), we address some basic concepts related to holographic quantum complexity. In section (5), we demonstrate that $H_{0}$ can be related to holographic complexity. Section (6) is devoted to conclusions and discussions.
@@ 2 The holographic principle as a prelude to the emergence of spacetime
A forthcoming quantum theory of gravity must provide a radical change in our comprehension of the world. However, possibly the most disruptive lesson quantum gravity must bring us is that related to the role of spacetime in our description of physical reality. In this sense, the holographic principle has been pointed as a possible guide [23, 24, 25].
@@
The most successful form of the holographic principle consists of the AdS/CFT correspondence which consists of a non-perturbative approach to superstring theory [26]. In its original formulation, the AdS/CFT correspondence establishes a duality between a type IIB string theory living in a $AdS_{5}\times S^{5}$ bulk and a conformal field theory (CFT) living on the bulk boundary.
@@
From the necessity of aligning the AdS/CFT ideas with the gravitational dynamics present in our four-dimensional universe, it has been considered its complementarity with the Randall-Sundrum II (RSII) scenario [27, 28, 29, 30]. By taking this way, our four-dimensional universe, which is viewed in the AdS/CFT framework as localized on the AdS boundary, turns to be described by a stack of coincident Randall-Sundrum II branes, and the CFT must be considered coupled to gravity.
Such a complementarity between the AdS/CFT and the Randall-Sundrum frameworks has given birth to a new version of the AdS/CFT correspondence, which has been proposed by Emparan, Kaloper, and Fabri [31], and independently by Tanaka [32] in 2002. It was named by Gregory as the Black Hole Holographic Conjecture (BHHC) [33, 34].
@@
This new version of the AdS/CFT correspondence proposes that the gravitational theory living in the AdS bulk, described by a type IIB string theory, must be holographically related not only to a CFT but to a semiclassical gravitational theory (gravity + quantum corrections) living on the AdS boundary. In such a context, bulk higher-dimensional effects must be translated as quantum corrections to the boundary gravitational theory.
@@
It has opened up the possibility of interesting investigations related to the nature of the boundary theory in the AdS/CFT framework. For example, by applying the BHHC ideas to a cosmological scenario, it has been demonstrated that quantum-corrected Friedmann equations similar to Loop Quantum Cosmology (LQC) ones can be induced on the AdS boundary [20].
@@
Due to the fact that quantum corrections arising in such a context must correspond to the holographic dual of higher-dimensional effects in the AdS bulk, such a result has raised the interesting possibility of connecting the superstring theory in the bulk to a quantum gravitational theory living on the AdS boundary.
@@
In fact, it has been demonstrated that LQG polymer quantization can be applied in such a context. In this case, the quantum degrees of freedom belonging to the boundary gravitational theory correspond to $U(1)$ quantum holonomies which give rise to quantum geometrical states $\psi(x)=e^{ipx}$ defined on the AdS boundary in the following way:
@@
Even though such holonomies are similar to that we have in LQG to describe a quantum spacetime, in the scenario introduced in [20], the quantization parameter defining such objects is related to the brane tension and, consequently, to the closed string coupling [35, 36].
In fact, the quantity $\Delta$ in Eq. (1) correspond to a quantum of area parameter given by [20]:
where $\sigma$ is the brane tension, $\alpha^{\prime}$ is the Regge slope, $g_{s}$ is the closed string coupling, $L$ is the AdS curvature radius, and $N$ corresponds to the number of branes sourcing gravitationally the bulk geometry [35, 36].
@@
It establishes a reinterpretation of branes as $U(1)$ quantum polymer structures. In this way, a possible conection between the main approaches to quantum gravity nowadays is traced, by connecting fundamental objects in string theory and LQG: branes and polymer structures.
Such a result has been useful in addressing interesting problems as those related to the big bang singularity in $AdS/CFT$ correspondence [37, 38], and the Immirzi ambiguity in LQG [39]. However, possible the main lesson it can bring us is that the way branes will be combined to source gravitationally the AdS geometry will depend on the rules governing quantum polymer structures similar to that appear in LQG.
However, at this point, some care needs to be taken since a decision must be made regarding the interpretation of using LQG techniques in the context of the AdS/CFT correspondence. In this sense, the existing LQG literature offers two possibilities.
@@
The first one is that the use of LQG techniques will provide the quantization of the AdS boundary geometry. However, it has been demonstrated that the LQG quantization techniques not only quantize spacetime but also modify it [40]. In this case, one should not have an AdS spacetime anymore after the application of the LQG quantization methods, and the bridge between LQG and string theory would be broken.
@@
The second, and most radical possibility, is that the LQG quantization procedure would not only quantize spacetime, but it will dissolve it into a pre-geometric structure that lives beyond spacetime itself, and from which spacetime ought to emerge.
@@
This is in tune with some AdS/CFT ideas that propose that spacetime must be an emergent aspect of physical reality. In fact, the AdS/CFT correspondence provides the most well-understood example for the emergence of spacetime by proposing that the AdS bulk must be emergent from the dynamics of quantum degrees of freedom living on the AdS boundary [41, 42, 43, 44, 45, 46]. However, the fundamental aspects of how the bulk spacetime, and its gravitational dynamics emerge from the dual theory, are still shrouded in many mysteries.
In this way, the theory introduced in [20] can bring us new insights on what seems to be in the heart of the $AdS/CFT$ correspondence: that spacetime must not be fundamental but ought to emerge from quantum information.
3 A new holographic prescription to the emergence of spacetime.
@@
The emergent nature of spacetime has been proved to be a peculiar feature of quantum gravity. In such a context, it has been proposed that spacetime must emerge from quantum information codified in a set of quantum reference frames connected by quantum correlations [47, 48, 49].
@@
The notion of quantum reference frame arises in the context of quantum mechanics as related to the universality of the theory: both observers and observed must satisfy the same natural laws. In this case, if the quantum description of the world is universal, one would expect that reference frames (observers) should obey the laws of quantum mechanics, corresponding in this way to quantum systems, such as quantum particles.
@@
Considering the string theoretical approach to quantum gravity, a brane is a physical object that generalizes the notion of a point particle. In this way, one can naturally suggest that branes must assume the role of quantum reference frames in such a context.
@@
Such a concern agrees with the result found out in [20] that branes correspond to $U(1)$ holonomies structures. In fact, this result allows us to conceive branes as quantum clocks, i.e., temporal quantum reference frames in a quantum gravity scenario.
@@
However, at this point, a peculiar detail related to the results found out in [20] arises: the holonomy structure of branes can be used not only to establish these objects as quantum reference frames but also to weave the entanglement network between them.
Such a conclusion is rooted in the fact that quantum holonomies $\hat{h}_{ij}$ consist of unitary transformations connecting two quantum geometrical states, belonging (in general) to two different Hilbert spaces, through entanglement [50, 51, 52]:
@@
In this way, if we label the $\mathcal{H}_{i}$s as the Hilbert spaces of $N$ branes, in the equations above, the indices $i,j$ will range as $i,j=1,...,N$, and the holonomies $\hat{h}_{ij}$ will be promoted to $U(N)$ unitary matrices.
However, as highlighted in [21], the important detail is that there is no fundamental distinction between the holonomies describing the branes serving as quantum reference frames and the holonomies establishing the quantum correlations between them. Note that, in such a construction, one could arbitrarily label some holonomies to correspond to that frozen to define the brane quantum reference frames, while taking the others to correspond to the connections among them. Consequently, even what has been called a quantum reference frame can be understood fundamentally as a quantum correlation, and vice-versa.
In this way, it is possible to introduce a structural realism [53] interpretation to quantum gravity in the context of the results found out in [20, 21], where only quantum correlations are considered as fundamental, and spacetime would be seen as completely emergent from the combinatorial rules governing the structure of such correlations.
Such a perspective can be useful to understand some of the most fundamental aspects related to the emergence of spacetime in quantum gravity. For example, according to the usual form of the holographic principle, spacetime geometry must emerge holographically from a quantum theory living in one spatial dimension lower. But a conceptual problem has been pointed out in such a construction: in the situation where the bulk emerges from the boundary, the last should be more fundamental than the first. It is difficult to reconcile with the fact that, in the AdS/CFT scenario, the duality between the boundary and the bulk theories is exact and thus symmetrical [54, 55, 56, 57].
This kind of conceptual barrier is rooted in the fact that it is a challenge to completely leave the idea of the fundamental existence of spacetime in physics. In fact, how could physics exist beyond spacetime, and how could things exist, and become entangled, for example, without some loci where and when they happen and change? In this case, in the context of the holographic principle, we still need the existence of a locus (the boundary) from which the bulk will emerge.
@@
It is possible to take a deeper look at the nature of the structure of quantum correlations from which spacetime must emerge by considering an interesting fact pointed out by Girelli and Livine in [47]. According to this fact, a vector in the representation space of the unitary group corresponds to an intertwiner with $N$ legs dressed by $SU(2)$ representations.
@@
Consequently, the $U(N)$ quantum states introduced in [21] can be described by a superposition of intertwiner states, and the geometry of the emergent AdS spacetime can be determined by the combinatorial rules governing abstract spin networks.
The results found out in [20, 21] make it possible to set out a new prescription for the emergence of spacetime in the context of the $AdS/CFT$ correspondence as follows:
A type IIB string theory living in a classical $AdS$ bulk emerges from abstract spin networks describing quantum correlations without correlata, in the limit where the number of correlations becomes large.
Important detail here is that abstract spin networks are graphs labeled by $SU(2)$ representations, in the same way we have in full LQG. However, such objects carry only combinatorial information, without any reference to a background geometry or topology. In this way, it is possible to say that, according to the scenario introduced in [21], spacetime must be seen as completely emergent from algebra.
We must highlight that the pre-geometrical regime from which spacetime ought to emerge must occur far from the limit where the number of branes sourcing the bulk geometry is large since, in such a limit, classical spacetime must emerge, according to the $AdS/CFT$ correspondence. It must correspond also to the limit where the number of quantum correlations becomes large.
Moreover, from the coincidence between the number of branes and the number of intertwiners external legs in the approach introduced in [21], the semiclassical limit proposed by the $AdS/CFT$ correspondence aligns with the reasoning introduced in [47] about the semiclassical limit of a pre-geometrical $U(N)$ matrix theory written in terms of $N$-leg intertwiners. In this context, the semiclassical limit is obtained when the number of intertwiners legs $N$ becomes large.
We note that in the context of the theory introduced in [20, 21], the concept of quantum complexity would be useful and could be seen as an intrinsic property of the structure of quantum correlations from which spacetime ought to emerge, being related to patterns of connectivity and transformation within such a structure.
In the next section, we shall explore the concept of holographic complexity in the context of the holographic principle. Moreover, we shall introduce the law governing its evolution.
@@
Briefly, complexity expresses the hardness of some task in terms of the amount of computational resources needed to accomplish it. In quantum information theory, any quantum state has an associated complexity, and preparing it using unitary operations is a task in itself. In this context, quantum entanglement has been pointed as the key resource to computational operations [62, 63].
@@
Considering the AdS/CFT formulation of the holographic principle, it has been suggested that the complexity associated with the quantum entanglement structure of conformal quantum fields on the boundary of an AdS space might give us insight into the emergence of the geometrical properties of the bulk. [64].
The origin of such an argument is the fact that entanglement alone is not enough to explain how complex bulk geometries emerge [65, 66]. However, by connecting the bulk spacetime properties to the complexity of quantum states, it is possible to trace an alternative way to understand how larger and more intricate bulk geometries can emerge.
Such an idea can be captured by the two main proposals to calculate quantum complexity in the context of the AdS/CFT correspondence, the so-called "Complexity equals to action" [67], and the "Complexity equals to volume conjectures [65, 66].
The first proposal says that the complexity of a quantum state on the AdS boundary must be related to the gravitational dynamics in the bulk spacetime, being proportional to the action in the Wheeler-DeWitt patch. The second proposal suggests that quantum complexity must be related to a volume in the AdS bulk, being proportional to the volume of the maximal codimension-1 bulk slice.
Concerning the second proposal, which will be implemented in the present letter, the holographic complexity $C_{h}$ related to some region on the AdS boundary is given by [64]:
@@
where $L$ is the AdS radius, $\gamma$ is the Ryu-Takayanagi (RT) surface [41, 42], and $Vol(\gamma)$ the volume under such a surface in the AdS bulk. The parameter $\epsilon$ denotes a necessary UV cutoff length scale in the boundary theory, and $Vol(R)$ denotes the volume of the boundary spacelike subregion, with a radial size $R$, which goes as $R^{d-1}$. Here $d$ gives the number of bulk spatial dimensions.
Considering the complexity measure above, if quantum complexity increases, bulk geometry also expands. It reinforces the idea that spacetime must be fundamentally linked to quantum computation, being not fundamental but emergent from quantum information.
An important aspect in such discussions is that quantum complexity must obey a second law analogous to the thermodynamical one for entropy [22, 68, 69]. According to such a law: If the computational complexity is less than the maximum, then with the overwhelming likelihood it will increase, till it saturates at a maximal value.
The second law of quantum complexity must have a crucial role in the issue of the emergence of spacetime since it must dictate the way spacetime must emerge from quantum information. In fact, even more implying that spacetime consists of an emergent entity, the second law of quantum complexity must imply that it must emerge in a deterministic way based on the rules of quantum computation.

### v4

@@ Abstract
In this paper, we propose a relationship between the so-called Hubble-Lemaître constant $H_{0}$ and the holographic complexity related to the emergence of spacetime in quantum gravity. Such a result can represent an important step to understanding the Hubble tension by introducing a quantum gravity perspective for cosmological observations, regarding the degree of quantum complexity we measure around us.
@@ 1 Introduction
In recent years, a controversy has arisen in cosmology, known as the Hubble tension [1, 2, 3]. This problem has its origin in a non-concordancy between early-universe measurements of the so-called Hubble-Lemaître constant $H_{0}$ provided by the Planck collaboration [4], and late-time observations provided by SHOES [5]. Such a discrepancy has been confirmed by several other late-time data [5, 6, 7, 8, 9, 12, 10, 11], which have shown a clear distinction from the value provided by the Planck satellite.
The early-universe result is based on the analysis of the Cosmic Microwave Background radiation (CMB), which was formed when the universe was in its childhood. On the other hand, late-time measurements are based on several observations related to objects formed in epochs when the universe was older and more complex, such as Cepheids, quasars, and red giant stars.
Putting it in numbers, the early universe result has provided a value of $H_{0}=67.4\pm 0.5\;km\;s^{-1}Mpc^{-1}$. On the other hand, the late-time universe measurements have provided values for $H_{0}\geq 70$. For example, the latest SHOES measurement gives $H_{0}=73.2\pm 1.3\;km\;s^{-1}Mpc^{-1}$ [11]. Moreover, several middle-time observations have provided intermediate values [13, 14].
@@
[…]
A forthcoming quantum version of cosmology, based on a still unknown theory of quantum gravity, can promote such a breakdown by reformulating our view about one of the most fundamental aspects of physical reality: spacetime itself. However, so far, no satisfactory proposal to solve the Hubble tension has been introduced by any candidate theory of quantum gravity.
In the present paper, we propose a possible explanation for the origin of the Hubble tension by considering new results on the nature of spacetime in quantum gravity introduced in [22, 23]. According to such results, spacetime must not be fundamental but emergent from quantum correlations without correlata, a reminiscent idea from Mermin’s interpretation of quantum mechanics [24, 25]. In such a context, we shall find that $H_{0}$ must be related to which must arise as an important element to understand the quantum aspect of the cosmos: the quantum complexity associated with the emergence of spacetime.
Consequently, the discrepancies between the $H_{0}$ values, when one considers the early and late-time measurements, can be attributed to a natural feature of the cosmos: the evolution of its complexity throughout time. Such a result may not only shed light on the Hubble tension but also introduce a new perspective for cosmological observations: in terms of holographic quantum complexity, we measure around us.
The paper is organized as follows: in sections (2) and (3), we review the main aspects of the theory introduced in [22, 23]. In section (4), we address some basic concepts related to holographic quantum complexity. In section (5), we demonstrate that $H_{0}$ can be related to holographic complexity required for the emergence of spacetime. In section (6), we show that the relationship between $H_{0}$ and holographic complexity can reproduce the results given for $H_{0}$ by early, middle, and late-time measurements. Section (8) is devoted to conclusions and discussions.
@@ 2 The holographic principle as a prelude to the emergence of spacetime
A forthcoming quantum theory of gravity must provide a radical change in our comprehension of the world. However, possibly the most disruptive lesson quantum gravity must bring us is that related to the role of spacetime itself in our description of physical reality.
In this sense, the holographic principle has been pointed out as a possible guide to quantum gravity [26, 27, 28]. According to such a principle, the gravitational dynamics in a region of spacetime can be completely described by a quantum theory living on its boundary.
@@
The most successful form of the holographic principle is the so-called AdS/CFT correspondence consisting of a non-perturbative approach to superstring theory [29]. In its original formulation, the AdS/CFT correspondence establishes a duality between a type IIB string theory living in a $AdS_{5}\times S^{5}$ bulk and a conformal field theory (CFT) living on the bulk boundary.
@@
From the necessity of aligning the AdS/CFT ideas with the gravitational dynamics present in our four-dimensional universe, it has been considered its complementarity with the Randall-Sundrum II (RSII) scenario [30, 31, 32, 33], another important insight in the string theoretical setup. By taking this way, our four-dimensional universe, viewed in the AdS/CFT framework as localized on the AdS boundary, turns out to be described by a stack of coincident Randall-Sundrum II branes, and the CFTs must be coupled to gravity.
@@
Such a complementarity between the AdS/CFT and the Randall-Sundrum frameworks has given birth to a new version of the AdS/CFT correspondence, which has been proposed by Emparan, Kaloper, and Fabri [34], and independently by Tanaka [35] in 2002. It was named by Gregory as the Black Hole Holographic Conjecture (BHHC) [36, 37], and proposes that the gravitational theory living in the AdS bulk, described by a type IIB string theory, must be holographically related not only to a CFT but to a semiclassical gravitational theory (gravity + quantum corrections) living on the AdS boundary. In such a context, bulk higher-dimensional effects must be translated as quantum corrections to the boundary gravitational theory.
@@
The BHHC has opened up the possibility of interesting investigations on the nature of the gravitational degrees of freedom belonging to the boundary theory in the AdS/CFT framework. For example, by applying its ideas to a cosmological scenario, it has been demonstrated that quantum-corrected Friedmann equations similar to Loop Quantum Cosmology (LQC) ones can be induced on the AdS boundary [22].
@@
Since quantum corrections arising in such a context must correspond to the holographic dual of higher-dimensional effects in the AdS space, such a result has raised the interesting possibility of connecting superstring theory living in the AdS bulk to a quantum gravitational theory living on the boundary.
@@
In fact, it has been demonstrated that LQG polymer quantization can be applied to the boundary gravitational theory in such a context. In this case, the quantum degrees of freedom belonging to this theory must correspond to $U(1)$ quantum holonomies giving rise to quantum geometrical states $\psi(x)=e^{ipx}$ defined on the AdS boundary in the following way:
@@
where $\sqrt{\Delta}$ consists of the lattice spacing parameter.
@@
It establishes a reinterpretation of branes as $U(1)$ quantum polymer structures, and a possible connection between the main approaches to quantum gravity nowadays can be traced by connecting fundamental objects in string theory and LQG: branes and polymer structures.
On the other hand, we note that even though the holonomies above are similar to that we have in LQG to describe a quantum spacetime, in the scenario introduced in [22], the quantization parameter defining such objects is related to the brane tension and, consequently, to the closed string coupling [38, 39]. In fact, the parameter $\Delta$ in Eq. (1) correspond to a quantum of area parameter given by [22]:
where $\sigma$ is the brane tension, $\alpha^{\prime}$ is the Regge slope, $g_{s}$ is the closed string coupling, $L$ is the AdS curvature radius, and $N$ corresponds to the number of branes sourcing gravitationally the AdS bulk geometry [38, 39]. We note, from Eq. (3), that the quantum of area $\Delta$ depends on the energy scale (or equivalently the length scale) used to probe spacetime geometry.
The results above have been useful in addressing interesting problems as those related to the big bang singularity in $AdS/CFT$ correspondence [40, 41], and the Immirzi ambiguity in LQG [42]. In this context, it has been also demonstrated that a positive cosmological constant can be induced on the AdS boundary, which has open up the possibility of explore possible observational evidences of the AdS/CFT correspondence.
Amidst all these results, the main lesson we have is that the way branes must be combined to source gravitationally the AdS geometry will depend on the rules governing quantum polymer structures, similar to those that appear in LQG. However, at this point, some care is needed in interpreting the use of LQG techniques in the context of the AdS/CFT correspondence. In this sense, the existing LQG literature offers two possibilities.
@@
The first one is that the use of LQG techniques will provide the quantization of the AdS boundary geometry. However, it has been demonstrated that the LQG quantization techniques not only quantize spacetime but also modify it [43]. In this case, one should not have an AdS spacetime anymore after the application of the LQG quantization methods to the AdS boundary, and the bridge between LQG and string theory would be broken.
@@
The second, and most radical possibility offered by the LQG literature, is that the LQG quantization procedure would not only quantize spacetime, but it would also dissolve it into a pre-geometric structure that lives beyond spacetime itself, and from which spacetime ought to emerge.
@@
Such a perspective is in tune with some AdS/CFT ideas that propose spacetime must be an emergent aspect of physical reality [44, 45, 46, 47, 48, 49]. In fact, the AdS/CFT framework provides the most well-understood example for spacetime emergence in quantum gravity by proposing that the AdS bulk must be emergent from quantum entanglement related to quantum degrees of freedom living on the AdS boundary.
However, the fundamental aspects of how bulk spacetime and its gravitational dynamics emerge from the dual theory remain shrouded in many mysteries. For example, it has been suggested that entanglement is not enough to explain the emergence of spacetime, but quantum complexity must also play a fundamental role in this process. As we shall see, the theory introduced in [22] may bring us new insights into this issue.
3 A holographic prescription to the emergence of spacetime in quantum gravity.
@@
In order to build a background-independent approach to quantum gravity, it has been proposed that spacetime must emerge from quantum information codified in a set of quantum reference frames connected by quantum correlations [50, 51, 52].
@@
The notion of a quantum reference frame arises in a quantum mechanical description of the world as related to the universality of the theory: both observers and observed must satisfy the same natural laws. In this case, if the quantum description of the world is universal, one would expect that reference frames (observers) should obey the laws of quantum mechanics, corresponding in this way to quantum systems, such as quantum particles.
@@
However, what kind of system could be considered as a quantum reference frame in the context of quantum gravity? Taking into account the string theoretical setup, a brane is a physical object that generalizes the notion of a point particle. In this way, one can naturally suggest that these objects can assume the role of quantum reference frames in a quantum gravitational context.
@@
By taking such a concern, the results found in [22] can provide the missing ingredient to conceive branes as quantum reference frames: that these objects can be seen as $U(1)$ quantum holonomy structures. In this case, one can see branes as quantum clocks, i.e., temporal quantum reference frames to the emergence of spacetime in quantum gravity [50, 51, 52].
@@
However, before proceeding, we note that a peculiar detail related to the scenario introduced in [22] arises in this discussion: the holonomic structure of branes can be used not only to establish these objects as quantum reference frames but also to weave the entanglement network between them. Such a conclusion is rooted in the fact that quantum holonomies $\hat{h}_{ij}$ consist of unitary transformations connecting two quantum geometrical states, belonging (in general) to two different Hilbert spaces, through entanglement [53, 54, 55]:
@@
In this context, if we label the $\mathcal{H}_{i}$s as the Hilbert spaces of $N$ branes, in the equations above, the indices $i,j$ will range as i,j=1,,Ni,j=1,...,N, and the holonomies $\hat{h}_{ij}$ will be promoted to $U(N)$ unitary matrices.
However, the important detail is that there is no fundamental distinction between the holonomies describing branes as quantum reference frames and the holonomies establishing the quantum correlations between them. Note that, in such a construction, one could arbitrarily label some holonomies to correspond to those frozen to define the brane quantum reference frames, while taking the others to correspond to the correlations between such objects. Consequently, even what has been called a quantum reference frame can be understood fundamentally as a quantum correlation, and vice versa.
In this case, as it was pointed out in [23], it is possible to take the universality of the theory one step further by introducing a scenario where there exists a unique kind of quantum object: quantum correlations. In this way, it is possible to introduce a Structural realistic interpretation to quantum gravity [56], where only quantum correlations are considered as fundamental, and quantum refence frames, as well as spacetime itself, would be seen as completely emergent from the combinatorial rules governing the structure of such correlations. Such ideas are reminiscent of the Mermin’s interpretation of quantum mechanics, where only quantum correlations must correspond to the elements of physical reality [24, 25].
Such a perspective, introduced in [23], can be useful to understand some of the most fundamental aspects related to the emergence of spacetime in quantum gravity. For example, according to the usual form of the holographic principle, spacetime geometry must emerge holographically from a quantum theory living in one spatial dimension lower. It corresponds to the usual AdS/CFT framework for the emergence of spacetime.
However, a conceptual problem has been pointed out in such a construction: in the situation where the bulk emerges from the boundary, the latter should be more fundamental than the former. It is difficult to reconcile with the fact that, in the AdS/CFT scenario, the duality between the boundary and the bulk theories is exact and thus symmetrical [57, 58, 59, 60].
This kind of conceptual barrier is rooted in the fact that it is a challenge to completely leave the idea of the fundamental existence of spacetime in physics. In fact, how could physics exist beyond spacetime, and how could things exist and become entangled, without some loci where and when they happen and change? In this case, in the usual form of the holographic principle, we still need the existence of a locus (the boundary) from which the bulk spacetime will emerge.
@@
By taking this route, we must focus on the combinatorial rules governing the structure of quantum correlations from which spacetime must emerge. To explore such rules, we have an interesting fact pointed out by Girelli and Livine [50] that a vector in the representation space of the unitary group corresponds to an intertwiner with $N$ legs dressed by $SU(2)$ representations.
@@
It introduces a very interesting description of the structure of quantum correlations from which spacetime must emerge, where the $U(N)$ quantum states describing such correlations can be written as a superposition of spin network states. In such a scenario, the geometry of the emergent AdS spacetime can be determined by the combinatorial rules governing spin networks.
An important detail here is that abstract spin networks are graphs labeled by $SU(2)$ representations, in the same way we have in full LQG. However, such objects carry only combinatorial information, without any reference to a background geometry or topology [50]. Consequently, it is possible to say that, according to the scenario introduced in [23], spacetime must be seen as completely emergent from the combinatorial rules governing abstract spin networks.
By considering such a scenario, the following prescription for the emergence of spacetime can be traced in the context of quantum gravity [22, 23]:
A type IIB string theory living in a classical $AdS$ bulk emerges from a superposition of abstract spin networks describing quantum correlations without correlate, in the limit where the number of correlations becomes large.
We highlight that the pre-geometrical regime from which spacetime ought to emerge must occur far from the limit where the number of branes sourcing the bulk geometry is large, since, in this regime, classical spacetime must emerge, according to the $AdS/CFT$ correspondence.
By observing that the number of branes sourcing gravitationaly the AdS spacetime coincides with the number of external legs of the intertwiners in the approach introduced in [23], we see that this reasoning is consistent with that of [50], which addresses the semiclassical limit of a pre-geometrical $U(N)$ matrix theory formulated in terms of $N$-leg intertwiners. In both cases, the semiclassical regime is reached as $N$, the number of intertwiner legs, becomes large. This large $N$ limit reflects a scenario where the number of quantum correlations also becomes large, matching the prescription discussed above.
In this way, in the theory proposed in [23], spacetime must be seen as emergent from quantum correlations. However, the way quantum correlations organize in order to give rise to spacetime emergence has not been proposed in this scenario.
In the next section, we shall explore the concept of holographic complexity in the context of the holographic principle, which can provide a possible route to the mechanism of spacetime emergence from quantum correlations: the second law of holographic complexity.
@@
Briefly, complexity expresses the hardness of some task in terms of the amount of computational resources needed to accomplish it. In quantum information theory, any quantum state has an associated complexity, and preparing it using unitary operations is a task in itself. In this context, quantum entanglement has been pointed out as the key resource to computational operations [65, 66].
@@
Considering the AdS/CFT formulation of the holographic principle, it has been suggested that the complexity associated with the quantum entanglement structure of conformal quantum fields on the boundary of an AdS space might give us insight into the emergence of the geometrical properties in the bulk. [67].
The origin of such an argument is the fact that entanglement alone is not enough to explain how complex bulk geometries emerge [68, 69]. However, by connecting the bulk spacetime properties to the complexity of quantum states, it is possible to trace an alternative way to understand how larger and more intricate bulk geometries can appear.
Such an idea can be captured by the two main proposals to calculate quantum complexity in the context of the AdS/CFT correspondence, the ”Complexity equals to volume[68, 69], and the ”Complexity equals to action[70] conjectures.
The first proposal suggests that quantum complexity must be related to a volume in the AdS bulk, being proportional to the volume of the maximal codimension-1 bulk slice. The second proposal says that the complexity of a quantum state on the AdS boundary must be related to the gravitational dynamics in the bulk spacetime, being proportional to the action in the Wheeler-DeWitt patch.
Concerning the ”Complexity equals to volume” proposal, the holographic complexity $C_{h}$ related to some region on the AdS boundary is given by [67]:
@@
where $L$ is the AdS radius, $\gamma$ is the Ryu-Takayanagi (RT) surface [44, 45], and $Vol(\gamma)$ is the volume under such a surface in the AdS bulk. The parameter $\epsilon$ denotes a necessary UV cutoff length scale in the boundary theory, and $Vol(R)$ denotes the volume of the boundary spacelike subregion, with a radial size $R$, which goes as $R^{d-1}$. Here $d$ gives the number of bulk spatial dimensions.
Considering the complexity measure above, if quantum complexity increases, bulk geometry also expands. In such a context, an important aspect is that quantum complexity must obey a second law analogous to the thermodynamic one for entropy. According to such a law, if the computational complexity is less than the maximum, then with overwhelming likelihood it will increase till it saturates at a maximal value [71, 72, 73].
The second law of quantum complexity must play a crucial role in the issue of the emergence of spacetime since it must dictate the way spacetime must emerge from quantum information. In the present context, it could tell us how spacetime emerges from quantum correlations. In fact, even more implying that spacetime consists of an emergent entity, the second law of quantum complexity implies that it must emerge in a deterministic way based on the rules of quantum computation.